Problem 1
Universal and BMT always dominate the milk powder market. That is because the two companies have vertical integration with milk raw material factories powder. Therefore, they have the competitive advantage of costly raw materials low. This has greatly helped the margins of the two companies. With margin very well, they have a strong position in the milk powder market.Universal and BMT plan to expand into the formula milk market. If they both entering the market, they can dominate the market with almost market share 100%.
Universal has ambitions to get a much higher market share than BMT. Currently, their competition is 'hot'. In fact, John Bryan, CEO Universal, won't want to be in the same room as Steven Voight, CEO of BMT. The feud between the two is quite great. You are the Marketing Manager of Universal. You are assigned to launch the formula milk. You can have 3 strategy options in launching products Formula milk:
1. Negotiate with BMT to agree to produce certain quantities
2. Competing 'head to head' by considering how BMJ will move (Non-Cooperative Strategy)?
3. Or encourage your company's R&D and production department to launch product earlier than BMT
You have done market research and know that market demand curve has the equation: P = 600 - Q. Assume your Marginal Cost is ‘0’ or zero .:
Question
1. What is the quantity (Q) in Cournot Equilibrium? How much will you suggest if you want to collude? What is the quantity (Q) in a competitive market?
2. Prepare Game Theory pay-off tables for every possibility? Show how is the calculation until you get the pay-off number.
3. Of the 3 strategies above, which one do you think is the best strategy for Universal and provide basic calculations and reasons why you chose the strategy. Consider various assumptions for all situations such as policy, investment, and regulation. Which of the 3 strategies does your boss like best? Give an explanation of your reasons.
4. The Pay-off table you made above is the Non-Dominant pay-off. Modify the pay-off table that you created in No-2 becomes the dominant pay-off strategy table. You can only modify the pay-off from BMT; Don't change the pay-off from Universal.
Answer #1

Answer:

Marginal revenue curve has twice the slope of aggregate demand curve and with the same vertical intercept.

Market demand = 600 - Q

MR = 600 - 2Q

Profit is maximised where MR = MC, MC=0

Therefore, 600 - 2Q = 0, Q = 300

1) From the above calculation, the Cournot Equilibrium Quantity is 300.In a collude situation the quantity should be such that price derived doesn't goes to zero. Say for example we expect that demand would increase by half, then Market demand curve will be P = 900 - 1.5Q. Quantity can be sold 500 units, for a Price P = 250(if we put Q = 500).

If both Universal and BMT cooperate, the quantity will be 300 units.

2)Situation Negotiate with BMT to agree to produce certain quantities Pay-off Matrix Universal BMT 500 3001 1 2 Competing head3) Of the 3 strategies above the best one is the 3rd option. Since, there is a cost which has been incurred in R&D, therefore Universal should take the advantage of this research and capitalise the increased demand as much as possible. My boss also liked the strategy.

4)  Situation Negotiate with BMT to agree to produce certain quantities Pay-off Matrix Universal BMT 500 500 1 2 Competing head

In the above table have changed the pay-off for quantity for BMT.

Situation 1: Negotiate, but still sell the same units as Universal.

Situation 3: BMT can always offer to sell more number of units than Universal, because Universal has incurred an R&D cost, which would form part of price to be charged to customer. Therefore, that portion of cost advantage BMT will have for a time duration, since all the R&D cost is recovered by Universal.